OpenStudy (anonymous):

elimination or Addition 1/2x+2/3y=1 1/4x-1/5y=-1/10

OpenStudy (anonymous):

What do you get post it here..

OpenStudy (anonymous):

@waterineyes I was confuse in messages. Lol my y=3/4

OpenStudy (anonymous):

y=6/8?

OpenStudy (anonymous):

Multiply first equation with 6 and second with 20 you will get: 3x + 4y = 6 5x - 4y = -2 Right??

OpenStudy (anonymous):

so I am going to use the equation where i use my LCM?

OpenStudy (anonymous):

Now just add them; 8x = 4 x = 1/2 Now put this value in any one equation: 3x + 4y = 6 Subtract 3x both the sides: 4y = 6-3x 4y = 6 - 3/2 4y = 9/2 y = 9/8

OpenStudy (anonymous):

ok. Thanks. I though i need to substitute x to the original equation

OpenStudy (anonymous):

You can substitute anywhere result will be the same..

OpenStudy (anonymous):

thanks @waterineyes becausei use 1/2(1/20+2/3y=1

OpenStudy (anonymous):

Suppose now I plug the values of x in second equation: 5x - 4y = -2 Subtract 5x: -4y = -2 - 5x -4y = -2 - 5/2 -4y = -9/2 multiply -1 both the sides: 4y = 9/2 y = 9/8..

OpenStudy (anonymous):

becausei use 1/2(1/2)+2/3y=1

OpenStudy (anonymous):

i go back to the first equation

OpenStudy (anonymous):

There you will have to use LCM and the solution becomes lengthy.. So you should put the calculated value in the expression which is easy to solve.

OpenStudy (anonymous):

thanks a lot again

OpenStudy (anonymous):

Welcome dear \(\huge \color{green} {\checkmark}\)

OpenStudy (anonymous):

the link that I read they always go back to the first equation though http://www.purplemath.com/modules/systlin5.htm

OpenStudy (anonymous):

In this case if we use very first equation then we have to take LCD and then we have to solve it becomes lengthy.. No doubt if your question is directly like this: 3x + 2y = 15 4x + 7y = 20 find x and y.. Here first you will find x and after that you will put it directly in first because this one is different from that fracitoned equation 1/2x + 1/2 y whatever it is.. So, try to keep things simple but not simpler..

OpenStudy (anonymous):

thanks again @waterineyes

OpenStudy (anonymous):

Welcome.. and you can note this point that all the equations given in that link are simple like: x + y = 2 3x + 3y = 6 Like this.. Am I right??

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Take care @nia_maepenix

OpenStudy (anonymous):

thanks for the help dear @waterineyes

OpenStudy (anonymous):

Happy To Help.. Welcome dear.. \(\huge \smile\)